A stone of mass 20g falls from rest from a height 8m calculate it's k.e after falling from a distance of 5m


1
Expert's answer
2021-01-26T17:50:15-0500

From the law of conservation of energy, we have:


"KE_{i,h=8\ m}+PE_{i,h=8\ m}=KE_{f,h=3\ m}+PE_{f,h=3\ m},""0+PE_{i,h=8\ m}=KE_{f,h=3\ m}+PE_{f,h=3\ m},""KE_{f,h=3\ m}=PE_{i,h=8\ m}-PE_{f,h=3\ m},""KE_{f,h=3\ m}=mg(h_f-h_i),""KE_{f,h=3\ m}=0.02\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot(8\ m-3\ m)=0.98\ J."